OpenStudy (anonymous):

"I NEED HELP WILL MEDAL AND FAN" what is the value of 3[29-(9+6)]-4*7? is the 4 a negative or a positive?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

3 [ 29 - (9+6) ] -4 * 7

Directrix (directrix):

Work from the inside out.

Directrix (directrix):

3* [ 29 - (9+6) ] -4 * 7 = 3 [ 29 - (15) ] -4 * 7

Directrix (directrix):

3 [ 29 - (15) ] -4 * 7 = 3 [ (14) ] -4 * 7 = Notice that the -4 has not come up yet.

OpenStudy (anonymous):

I found out the answer for that

Directrix (directrix):

3 [ (14) ] -4 * 7 = 3 [ (14) ] - 28 =

Directrix (directrix):

42 - 28 = 14 Is that what you got?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I posted a new question